OpenStudy (anonymous):

How do i integrate: Integral x^3sqrt(x^2+7)

OpenStudy (anonymous):

\[\int\limits_{}^{}x^3\sqrt{x^2+7}\]

OpenStudy (chaise):

Change the square root symbol to power half and integrate like normal. Do you need help with integrating after this step? x^2*(x^2+7)^(1/2)

OpenStudy (anonymous):

the integral is x^3sqrt(x^2+7)

OpenStudy (anonymous):

yes i still need help

OpenStudy (chaise):

Square root and to the power half is the same thing - try it on your calculator. \[x ^{3}\sqrt{x^2+7} = x^3 (x^2+7)^{1/2}\] \[\int\limits_{ }^{ }x^3 (x^2+7)^{1/2}=(1/15)(x ^{2}+7)^{3/2}(3x^2-14) +c\]

OpenStudy (anonymous):

you just game the answer, not how to do it

OpenStudy (anonymous):

what should i let u be to solve this?

OpenStudy (chaise):

(x^2 + 7)^(1/2)

OpenStudy (anonymous):

that would serve no purpose as du/dx wuld then be: 1/2(x^2+7)^-1/2 *(2x)

OpenStudy (chaise):

Then I guess your only option is x^3. ;)

OpenStudy (anonymous):

you dont know how to solve this do you

OpenStudy (anonymous):

hey zip..i haven't worked this out completely...but i would thinnk let u = x^2 +7

OpenStudy (anonymous):

yeah, that seems reasonable

OpenStudy (chaise):

I know perfectly well how to integrate by parts - but you seem like a master.

OpenStudy (anonymous):

du would then be 1/2 du = xdx

OpenStudy (anonymous):

come on dont pretend chaise

OpenStudy (anonymous):

right,that would get rid of one of the x's, but then i think i might need to rewrite x in terms of u

OpenStudy (anonymous):

i think you'll have to integrate a 2nd time b/c you'll have an x^2 left over

OpenStudy (anonymous):

again...this is completely by observation...i've not worked anythign out

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

help please

OpenStudy (nikvist):

\[\int x^3\sqrt{x^2+7}\,dx\quad,\quad y=x^2+7\quad,\quad dy=2xdx\] \[\frac{1}{2}\int(y-7)\sqrt{y}\,dy=\frac{1}{2}\int(y^{3/2}-7y^{1/2})\,dy=\frac{1}{5}y^{5/2}-\frac{7}{3}y^{3/2}+C=\] \[=y^{3/2}\left(\frac{1}{5}y-\frac{7}{3}\right)+C=\frac{1}{15}y^{3/2}\left(3y-35\right)+C=\] \[\frac{1}{15}(x^2+7)^{3/2}\left(3(x^2+7)-35\right)+C=\frac{1}{15}(x^2+7)^{3/2}\left(3x^2-14\right)+C\]